On Saturday, a local hamburger shop sold a combined total of 324 hamburgers and cheeseburgers. The number of cheeseburgers sold was three times the number of hamburgers sold. How many hamburgers were sold on Saturday?
Answer by ikleyn(52925)

.

You write an equation as you read the problem


    h + 3h = 324


Then you simplify it and find the answer


    4h     = 324

     h     = 324/4 = 81.


ANSWER.  81 hamburgers.

Solved.

Is something that you do not understand in my post ?